# Continual Lie algebras determined by chain complexes

### v4

@@ 1. Introduction
Continual Lie algebras introduced in [16] are generalizations of infinite-dimensional Lie algebras with systems of discrete roots [12]. Then the notion was generalized and developed in [17][20]. These algebras are formulated in such a way that the space of roots is defined by continual sets of vectors. In commutation relations, the generators depend on kernels which are functionals of continual roots. Jacobi identity for continual Lie algebras result in non-trivial functional relations for kernels. Similar constructions appeared recently in the study of certain Hall algebras of coherent sheaves described in terms of certain continuum limits of Kac-Moody algebras [1].
@@
Though the general theory of continual Lie algebras and their representations is missing, there exist many applications, especially in the theory of completely integrable and exactly solvable models [14]. Various applications can be also found in other fields of mathematics [4, 6, 20].
@@
In original papers [16][20], in order to find new classes of examples of continual Lie algebras various approaches were studied. In a series of papers [18, 19] the authors have generalized first known examples and have found new non-trivial and fundamental ones arising from various branches of mathematics, in particular, functional analysis [20], differential geometry, algebraic geometry, non-commutative geometry, and mathematical physics [4, 8]. Some of them came from physical applications, e.g., equations related to the Ricci flow and cosmology [4, 8]. Though they appear in various corners of modern mathematical research, at some stage, the arsenal of ideas of generation of new examples of continual Lie algebras was exhausted. An intriguing question is how to construct a general way which would allow to create continual Lie algebras. In many cases, this problem reduces to the question of finding new appropriate solutions for relations (1.1) (which is a very interesting problem by itself) following from Jacobi identity.
@@
The idea of this work is to use properties of general chain complexes, (with complex spaces depending on sets of parameters) in order to derive the structure of continual Lie algebras. The plan of the paper is the following. In the next subsection we recall the notion of a continual Lie algebra [16]. In Section 2, starting from a general (infinite) chain complex, we first define a product acting on elements of various spaces of a complex. We require the Leibniz rule to be fulfilled for (co)boundary operators with respect to this product. In addition to that we assume its skew-symmetry properties. Thus we endow the complex with the structure of a graded differential algebra $\mathcal{C}$.
@@
We then apply the natural condition of orthogonality with respect to a product among the chain complex spaces. For instance, in differential geometry, the orthogonality condition reduces to the integrability condition for differential forms. Further actions of (co)boundary operators, as well as consequences from the definition of the multiplication, generate a system of differential relations for elements of various chain complex spaces with extra compatibility conditions for indices. This determines the structure of a graded algebra with differential relations for $\mathcal{C}$. By choosing independent elements satisfying the above relations, we then show the main result of this paper: for a graded differential algebra associated to a chain complex, the independent generators, the system of differential relations, and Jacobi identity define the structure of a continual Lie algebra with the root space provided by the spaces parameters of a chain complex.
@@
In Section 3 we consider the example of the Čech-de Rham cochain bicomplex [5] (c.f. the Appendix) associated to a smooth manifold foliation [7, 2, 3]. This bicomplex has a deep geometric meaning [9] and is defined for the spaces of differential forms and holonomy embeddings acting between sections of the transversal basis for a foliation. In the Appendix we recall the notions of holonomy mappings, transversal sections, and transversal basis for a foliation. The space of differential forms has already a structure of bigraded differential algebra with respect to the natural product [5]. For the bicomplex under consideration, we define the second product, satisfying properties required for the construction of a system of differential relations. According to the exposition we give in Section 2, by involving the orthogonality conditions for an arbitrary pairs of the Čech-de Rham bicomplex, we apply the general scheme of construction of a bigraded differential algebra. Then we single out generators and commutation relations for the corresponding continual Lie algebra with the space of roots and kernels given by the sets of holonomy embeddings. In a particular geometric case associated to a codimension one foliation over a three-dimensional smooth manifold, we start from the integrability condition applied to elements of the same bicomplex space which leads to the appearance of a continual Lie algebra with generators, kernels, and commutation relations explicitly described.
@@ 1.1. Continual Lie algebras
for all $\phi$, $\psi\in E$. It is also assumed that Jacobi identity is satisfied. Then the conditions on mappings $\widehat{K}=\left(K_{0,0},K_{0},K_{\pm 1}\right)$ follow:
@@
for all $\phi$, $\psi$, $\chi\in E$. An infinite dimensional algebra $\mathcal{G}\left(E;\widehat{K}\right)=\mathcal{G}^{\prime}\left(E;\widehat{K}\right)/J$, is called a continual contragredient Lie algebra, where $\mathcal{G}^{\prime}(E;\widehat{K})$ is a Lie algebra freely generated by ${\widehat{\mathcal{G}}}$, and $J$ is the largest homogeneous ideal with trivial intersection with $\mathcal{G}_{0}$ (consideration of the quotient is equivalent to imposing the Serre relations in the case of finite-dimensional simple complex Lie algebras) [18, 19]. $\mathcal{G}$ is endowed with a $\mathbb{Z}$-grading $\mathcal{G}=\bigoplus_{n\in\mathbb{Z}}\mathcal{G}_{n}$, where elements of subspaces $\mathcal{G}_{n}$ satisfy the standard grading condition $\left[\mathcal{G}_{n},\mathcal{G}_{m}\right]\subset\mathcal{G}_{n+m}$, where $X_{n}(\phi)\in\mathcal{G}_{n}$, and higher mapping $K_{n,m}(\phi,\psi)$ are present in commutation relations among $X_{n}(\phi)$ and $X_{n}(\psi)$. In general, Jacobi identity for generators that belong to all grading spaces has the form
@@ 2.1. The graded differential algebra
We assume also that the chain complex above has some properties of a algebra with respect to the product (2.2). Namely, consider the elements $\Phi_{k}\in C(k,\Theta_{k})$, and $\Phi_{l}\in C(l,\Theta_{l})$. Let us assume, in particular, that the product (2.2) satisfies the condition
@@ 2.3. Construction of a continual Lie algebra from a chain complex
where brackets mean the ordinary commutator. It is known that the introduction of the commutator with respect to the original multiplication of an algebra transfers it into a Lie algebra when Jacobi conditions are satisfied. In our case we show that the differential algebra $\mathcal{C}$ defined above being supplied with the orthogonality conditions deliver the structure of a continual Lie algebra,
The setup of this subsection combined with the previous subsection provides us with a proof of the main result of this paper:
Proposition 1.
For the set $I_{0}$ of all pairs of independent elements $\alpha_{i}^{(K_{i})}$, the orthogonality condition (2.5), the generators $\left\{\chi,\;\delta(n_{0})\chi,\;\Phi,\;\delta(n)\Phi,\;\alpha_{i}^{(K_{i})},\;\delta(n(i))\alpha_{i}^{(K_{i})}\right\}$, $n(i)=n_{i}^{(K_{i})}$, the relations (2.2), and Jacobi identity (1.1) form a continual Lie algebra $\mathcal{G}(\Theta_{n(i)})$ with the root space depending on the set of parameters $\Theta_{n(i)}$, of spaces $C\left(n(i),\Theta_{n(i)}\right)$.
Though the structure of the system (2.2) may seem to be not very complicated, actual properties of the corresponding continual Lie algebra depends on properties of the spaces of a specific the bicomplex (2.1) and the nature of parameters $\Theta_{i}$. For a fixed choice of a path of independent functions/differential equations in the system (2.2), there exists a variety of choices on how to identify generators of a continual Lie algebra with generators of the differential algebra. Therefore, the actual form of commutation relations for the corresponding continual Lie algebra varies accordingly. One can also chose various ways how to define a grading for each specific $\mathcal{G}(\Theta_{n(i)})$, for generators of a continual Lie algebra (see subsection 1.1) resulting from (2.2). The structure of the product (2.2) together with the condition (2.3), and the action of the differentials $\delta$ provide Jacobi identity for generators on the continual Lie algebra $\mathcal{G}(\Theta_{n(i)})$, and, simultaneously, apply conditions of the form (1.1) to elements of the parameter spaces $\Theta_{(n(i))}$, $i\geq 0$. In the next Section we specify the above construction and Proposition (1) in the case of double cochain complex (5.1)–(5.2) associated with differential forms [5]. We derive explicitly the generators and commutation relations for the corresponding continual Lie algebras.
@@ 3. An example: double complex associated with foliations
Recall the formulation of the Čech-de Rham cohomology given in [5] for foliations on smooth manifolds (see the Appendix). In this case the spaces in (2.1) are differential forms $C^{n,m}(\mathcal{F})$ defined on a foliation, and the coboundary operators is given by $\delta^{p,q}=(-1)^{p}d+\delta^{q}$, where $d$ and $\delta$ operators are defined in the Appendix. The ordinary product for differential forms $\omega^{n,m}(h_{1},\ldots,h_{m})\in C^{n,m}\left(\mathcal{F}\right)$ is given by (5.3). The product (2.2) required for the formulation of Section 2 is provided by the commutator (2.17). As it was mentioned in [5], the bicomplex (5.1)–(5.2) has the structure of a bigraded differential algebra with respect to the ordinary product (5.3). According to explanations of Section 2, for the bicomplex (5.1)–(5.2) generate the graded differential algebra with relations (2.2) with respect to the product (2.17).
@@
Recall that according to Theorem 1 of [5], the definition of the spaces $C^{n,m}$ do not depend on the choice of the transversal basis $\mathcal{U}$, still it depends on parameters of foliation $\mathcal{F}$. Nevertheless, differential forms $\omega^{n,m}$ do depend on holonomy mappings (see the Appendix) $h_{j}$, $j\geq 0$, and play the role of extra parameters in the consideration.
As explained at the end of Section 2, due to Proposition (1), a path $(K_{i})$, $i\geq 1$, defining the generators $\left\{\chi,\;\delta^{n_{0},m_{0}}\chi,\;\Phi,\;\delta^{n,m}\Phi,\;\alpha_{i}^{(K_{i})},\;\delta^{n(i),m(i)}\alpha_{i}^{(K_{i})}\right\}$, $n(i)=n_{i}^{(K_{i})}$, $m(i)=m_{i}^{(K_{j})}$, and relations for independent elements of $C^{n,m}$ in (2.2), form a continual Lie algebra $\mathcal{G}(\mathcal{F})$ with the space of roots provided by the holonomy mappings and of parameters of the foliation $\mathcal{F}$.
@@ 3.1. Double cochain complex: Godbillon-Vey type example
In this subsection we provide the explicit example for Proposition (1) in the case of the orthogonality condition (2.5) applied to the particular case when, in the consideration of the previous subsection, $\Phi=\chi$. In particular, in differential geometry, the case of a foliation $\mathcal{F}$ of codimension one defined by a one-form on a three-dimensional manifold, and the formulation of the Godbillon-Vey cohomology class, are included in this consideration.
@@
Let us denote ${\bf h}_{n}=\left(h_{1},\ldots,h_{n}\right)$, an $n$-tuple of holonomy mappings (see the Appendix). In this case, for forms $\left\{\chi,\;\delta^{n,m}\chi,\;\alpha^{(R)}_{1},\;\delta^{r+1,t+1}\alpha_{1}^{(L)}\right\}$, define the following continual Lie algebra by identifying the differential forms with generators of $\mathcal{G}(\mathcal{F})$ as
@@
Note that an element of $C^{n,m}(\mathcal{F})$ is an $n$-form $\omega(h_{1},\ldots,h_{n})$ depending on $n$ holonomy maps. Thus the space of continual roots is provided by the space of holonomy embeddings (see the Appendix). Taking into account (3.1), we find the kernels
@@
It is easy to check that Jacobi identity for generators (3.8) are fulfilled. In the case of codimension one foliation defines by a one-form, $\chi$, $\alpha^{(R)}_{1}$, $\chi\in C^{1,m}(\mathcal{F})$. Recall [10] that the Godbillon-Vey cohomology class is given by $\left[\alpha^{(R)}_{1}\wedge\delta^{1,m}\alpha^{(R)}_{1}\right]$. The construction above clarifies the Lie-algebraic meaning of this cohomology class.

### v5

@@ 1. Introduction
Continual Lie algebras introduced in [16] are generalizations of infinite-dimensional Lie algebras with systems of discrete roots [12]. Then the notion was generalized and developed in [17][20]. These algebras are formulated in such a way that the space of roots is defined by continual sets of vectors. In commutation relations, the generators depend on kernels which are functionals of continual roots. the Jacobi identity for continual Lie algebras result in non-trivial functional relations for kernels. Similar constructions appeared recently in the study of certain Hall algebras of coherent sheaves described in terms of certain continuum limits of Kac-Moody algebras [1].
@@
Though a general theory of continual Lie algebras and their representations is missing, there exist many applications, especially in the theory of completely integrable and exactly solvable models [14]. Various applications can be also found in other fields of mathematics [4, 6, 20].
@@
In original papers [16][20], in order to find new classes of examples of continual Lie algebras various approaches were studied. In a series of papers [18, 19] the authors have generalized first known examples and have found new non-trivial and fundamental ones arising from various branches of mathematics, in particular, functional analysis [20], differential geometry, algebraic geometry, non-commutative geometry, and mathematical physics [4, 8]. Some of them came from physical applications, e.g., equations related to the Ricci flow and cosmology [4, 8]. Though they appear in various corners of modern mathematical research, at some stage, the arsenal of ideas of generation of new examples of continual Lie algebras was exhausted. An intriguing question is how to construct a general way which would allow to create continual Lie algebras. In many cases, this problem reduces to the question of finding new appropriate solutions for relations (1.1) (which is a very interesting problem by itself) following from the Jacobi identity.
The idea of this work is to use properties of general chain complexes, (with complex spaces depending on sets of parameters) in order to derive the structure of continual Lie algebras. The main result of this paper consists in the following
Theorem 1.
For the set $I_{0}$ of all pairs of independent elements $\alpha_{i}^{(K_{i})}$, the orthogonality condition (2.5), the generators $\left\{\chi,\;\delta(n_{0})\chi,\;\Phi,\;\delta(n)\Phi,\;\alpha_{i}^{(K_{i})},\;\delta(n(i))\alpha_{i}^{(K_{i})}\right\}$, $n(i)=n_{i}^{(K_{i})}$, the relations (2.2), and the Jacobi identity (1.1) form a continual Lie algebra $\mathcal{G}(\Theta_{n(i)})$ with the root space depending on the set of parameters $\Theta_{n(i)}$, of spaces $C\left(n(i),\Theta_{n(i)}\right)$.
@@
The plan of the paper is the following. In the next subsection we recall the notion of a continual Lie algebra [16]. In Section 2, starting from a general (infinite) chain complex, we first define a product acting on elements of various spaces of a complex. We require the Leibniz rule to be fulfilled for (co)boundary operators with respect to this product. In addition to that we assume its skew-symmetry properties. Thus we endow the complex with the structure of a graded differential algebra $\mathcal{C}$.
@@
We then apply the natural condition of orthogonality with respect to a product among the chain complex spaces. For instance, in differential geometry, the orthogonality condition reduces to the integrability condition for differential forms. Further actions of (co)boundary operators, as well as consequences from the definition of the multiplication, generate a system of differential relations for elements of various chain complex spaces with extra compatibility conditions for indices. This determines the structure of a graded algebra with differential relations for $\mathcal{C}$. By choosing independent elements satisfying the above relations, we then show the main result of this paper: for a graded differential algebra associated to a chain complex, the independent generators, the system of differential relations, and the Jacobi identity define the structure of a continual Lie algebra with the root space provided by the spaces parameters of a chain complex.
@@
In Section 3 we consider the example of the Čech-de Rham cochain bicomplex [5] (c.f. subsection 3.1) associated to a smooth manifold foliation [7, 2, 3]. This bicomplex has a deep geometric meaning [9] and is defined for the spaces of differential forms and holonomy embeddings acting between sections of the transversal basis for a foliation. In subsection 3.1 we recall the notions of holonomy mappings, transversal sections, and transversal basis for a foliation. The space of differential forms has already a structure of bigraded differential algebra with respect to the natural product [5]. For the bicomplex under consideration, we define the second product, satisfying properties required for the construction of a system of differential relations. According to the exposition we give in Section 2, by involving the orthogonality conditions for an arbitrary pairs of the Čech-de Rham bicomplex, we apply the general scheme of construction of a bigraded differential algebra. Then we single out generators and commutation relations for the corresponding continual Lie algebra with the space of roots and kernels given by the sets of holonomy embeddings. In a particular geometric case associated to a codimension one foliation over a three-dimensional smooth manifold, we start from the integrability condition applied to elements of the same bicomplex space which leads to the appearance of a continual Lie algebra with generators, kernels, and commutation relations explicitly described.
@@ 1.1. Continual Lie algebras
for all $\phi$, $\psi\in E$. It is also assumed that the Jacobi identity is satisfied. Then the conditions on mappings $\widehat{K}=\left(K_{0,0},K_{0},K_{\pm 1}\right)$ follow:
@@
for all $\phi$, $\psi$, $\chi\in E$. An infinite dimensional algebra $\mathcal{G}\left(E;\widehat{K}\right)=\mathcal{G}^{\prime}\left(E;\widehat{K}\right)/J$, is called a continual contragredient Lie algebra, where $\mathcal{G}^{\prime}(E;\widehat{K})$ is a Lie algebra freely generated by ${\widehat{\mathcal{G}}}$, and $J$ is the largest homogeneous ideal with trivial intersection with $\mathcal{G}_{0}$ (consideration of the quotient is equivalent to imposing the Serre relations in the case of finite-dimensional simple complex Lie algebras) [18, 19]. $\mathcal{G}$ is endowed with a $\mathbb{Z}$-grading $\mathcal{G}=\bigoplus_{n\in\mathbb{Z}}\mathcal{G}_{n}$, where elements of subspaces $\mathcal{G}_{n}$ satisfy the standard grading condition $\left[\mathcal{G}_{n},\mathcal{G}_{m}\right]\subset\mathcal{G}_{n+m}$, where $X_{n}(\phi)\in\mathcal{G}_{n}$, and higher mapping $K_{n,m}(\phi,\psi)$ are present in commutation relations among $X_{n}(\phi)$ and $X_{n}(\psi)$. In general, the Jacobi identity for generators that belong to all grading spaces has the form
@@ 2.1. The graded differential algebra
We assume also that the chain complex above has some properties of an algebra with respect to the product (2.2). Namely, consider the elements $\Phi_{k}\in C(k,\Theta_{k})$, and $\Phi_{l}\in C(l,\Theta_{l})$. Let us assume, in particular, that the product (2.2) satisfies the condition
@@ 2.3. Construction of a continual Lie algebra from a chain complex
where brackets mean the ordinary commutator. It is known that the introduction of the commutator with respect to the original multiplication of an algebra transfers it into a Lie algebra when the Jacobi conditions are satisfied. In our case we show that the differential algebra $\mathcal{C}$ defined above being supplied with the orthogonality conditions deliver the structure of a continual Lie algebra.
The setup of this subsection combined with the previous subsection provides us with a proof of the main result of this paper, Theorem 1.
Though the structure of the system (2.2) may seem to be not very complicated, actual properties of the corresponding continual Lie algebra depends on properties of the spaces of a specific the bicomplex (2.1) and the nature of parameters $\Theta_{i}$. For a fixed choice of a path of independent functions/differential equations in the system (2.2), there exists a variety of choices on how to identify generators of a continual Lie algebra with generators of the differential algebra. Therefore, the actual form of commutation relations for the corresponding continual Lie algebra varies accordingly. One can also chose various ways how to define a grading for each specific $\mathcal{G}(\Theta_{n(i)})$, for generators of a continual Lie algebra (see subsection 1.1) resulting from (2.2). The structure of the product (2.2), the condition (2.3), and the action of the differentials $\delta$ provide the Jacobi identity for generators on the continual Lie algebra $\mathcal{G}(\Theta_{n(i)})$, and, simultaneously, apply conditions of the form (1.1) to elements of the parameter spaces $\Theta_{(n(i))}$, $i\geq 0$. In the next Section we specify the above construction and Theorem (1) in the case of double cochain complex (3.1)–(3.2) associated with differential forms [5]. We derive explicitly the generators and commutation relations for the corresponding continual Lie algebras.
@@ 3. An example: double complex associated with foliations
3.1. Čech-de Rham complex for foliations
In this subsection we recall [5] the notion of the basis of transversal sections, and Čech-de Rham complex for a foliation of a smooth manifold. Let $M$ be a smooth manifold of dimension $n$, equipped with a foliation $\mathcal{F}$ of co-dimension $l$ [5]. A transversal section of $\mathcal{F}$ is an embedded $l$-dimensional submanifold $U\subset M$ which is everywhere transverse to $\mathcal{F}$ leaves. If $\alpha$ is a path between two points $x$ and $y$ on the same leaf, and if $U$ and $V$ are transversal sections through $x$ and $y$, then $\alpha$ defines a transport along the leaves from a neighborhood of $x$ in $U$ to a neighborhood of $y$ in $V$. Therefore, we define a germ of a diffeomorphism $hol(\alpha):(U,x)\longrightarrow(V,y)$, called the holonomy of the path $\alpha$. If such a transport is defined in all of $U$ and embeds $U$ into $V$, this embedding $h:U\hookrightarrow V$ is sometimes also denoted by $hol(\alpha):U\hookrightarrow V$. Embeddings of this form will be called holonomy embeddings.
The product is defined by
3.2. The double complex
In the previous subsection we recalled the formulation of the Čech-de Rham cohomology given in [5] for foliations on smooth manifolds (see subsection 3.1). In this case the spaces in (2.1) are differential forms $C^{n,m}(\mathcal{F})$ defined on a foliation, and the coboundary operators is given by $\delta^{p,q}=(-1)^{p}d+\delta^{q}$, where $d$ and $\delta$ operators are defined in subsection 3.1. The ordinary product for differential forms $\omega^{n,m}(h_{1},\ldots,h_{m})\in C^{n,m}\left(\mathcal{F}\right)$ is given by (3.3). The product (2.2) required for the formulation of Section 2 is provided by the commutator (2.17). As it was mentioned in [5], the bicomplex (3.1)–(3.2) has the structure of a bigraded differential algebra with respect to the ordinary product (3.3). According to explanations of Section 2, for the bicomplex (3.1)–(3.2) generate the graded differential algebra with relations (2.2) with respect to the product (2.17).
@@ 3.2. The double complex
Recall that according to Theorem 1 of [5], the definition of the spaces $C^{n,m}$ do not depend on the choice of the transversal basis $\mathcal{U}$, still it depends on parameters of foliation $\mathcal{F}$. Nevertheless, differential forms $\omega^{n,m}$ do depend on holonomy mappings (see subsectioncm) $h_{j}$, $j\geq 0$, and play the role of extra parameters in the consideration.
As explained at the end of Section 2, due to Theorem (1), a path $(K_{i})$, $i\geq 1$, defining the generators $\left\{\chi,\;\delta^{n_{0},m_{0}}\chi,\;\Phi,\;\delta^{n,m}\Phi,\;\alpha_{i}^{(K_{i})},\;\delta^{n(i),m(i)}\alpha_{i}^{(K_{i})}\right\}$, $n(i)=n_{i}^{(K_{i})}$, $m(i)=m_{i}^{(K_{j})}$, and relations for independent elements of $C^{n,m}$ in (2.2), form a continual Lie algebra $\mathcal{G}(\mathcal{F})$ with the space of roots provided by the holonomy mappings and of parameters of the foliation $\mathcal{F}$.
@@ 3.3. Double cochain complex: Godbillon-Vey type example
In this subsection we provide the explicit example for Theorem (1) in the case of the orthogonality condition (2.5) applied to the particular case when, in the consideration of the previous subsection, $\Phi=\chi$. In particular, in differential geometry, the case of a foliation $\mathcal{F}$ of codimension one defined by a one-form on a three-dimensional manifold, and the formulation of the Godbillon-Vey cohomology class, are included in this consideration.
@@
Let us denote ${\bf h}_{n}=\left(h_{1},\ldots,h_{n}\right)$, an $n$-tuple of holonomy mappings (see subsection 3.1). In this case, for forms $\left\{\chi,\;\delta^{n,m}\chi,\;\alpha^{(R)}_{1},\;\delta^{r+1,t+1}\alpha_{1}^{(L)}\right\}$, define the following continual Lie algebra by identifying the differential forms with generators of $\mathcal{G}(\mathcal{F})$ as
@@
Note that an element of $C^{n,m}(\mathcal{F})$ is an $n$-form $\omega(h_{1},\ldots,h_{n})$ depending on $n$ holonomy maps. Thus the space of continual roots is provided by the space of holonomy embeddings (see subsection 3.1). Taking into account (3.3), we find the kernels
@@
It is easy to check that the Jacobi identity for generators (3.11) are fulfilled. In the case of a codimension one foliation defined by a one-form, $\chi$, $\alpha^{(R)}_{1}$, $\chi\in C^{1,m}(\mathcal{F})$. Recall [10] that the Godbillon-Vey cohomology class is given by $\left[\alpha^{(R)}_{1}\wedge\delta^{1,m}\alpha^{(R)}_{1}\right]$. The construction above clarifies the Lie-algebraic meaning of this cohomology class.